OpenStudy (anonymous):

which numbers betwee 40 and 49 have 2 and 3 in their prime factorization? use exponents to write the factorization of each number

OpenStudy (anonymous):

if they have 2 and 3 they also have 6, so this question is asking what numbers between 40 and 49 are multiples of 6

OpenStudy (anonymous):

\[6\times 7=42\]works as does \[6\times 8=48\]

OpenStudy (anonymous):

thank you!!! this is now making sense

OpenStudy (anonymous):

how do i use exponents to write the factorization of each number?

OpenStudy (anonymous):

From a Mathematica one line program and a coder's text editor: {{{2, 3}, {5, 1}}, {{41, 1}}, {{2, 1}, {3, 1}, {7, 1}}, {{43,1}}, {{2, 2}, {11, 1}}, {{3, 2}, {5, 1}}, {{2, 1}, {23, 1}}, {{47,1}}, {{2, 4}, {3, 1}}, {{7, 2}} } 40=2^3 5^1 41=41^1 42=2^1 3^1 7^1 43=43^1 . . . . . 48=2^4 3^1 49=7^2